Circular No. 6533

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970111

E. Costa, M. Feroci, and L. Piro, Istituto Astrofisica Spaziale, C.N.R., Frascati; F. Frontera and G. Zavattini, Universitá di Ferrara; L. Nicastro and E. Palazzi, Istituto Tecnologie e Studio Radiazioni Extraterrestri, C.N.R., Bologna; G. Gandolfi, M. Smith, A. Coletta, G. D’Andreta, and G. Spoliti, BeppoSAX Science Operation Center, Roma; and R. Jager and J. Heise, Space Research Organization of the Netherlands, Utrecht, communicate: “The BeppoSAX Gamma Ray Burst Monitor (GRBM) was triggered on Jan. 11.405556 UT. Quick-look analysis carried out at the BeppoSAX Science Operation Center shows the following. From the GRBM lightcurve, we see an intense γ-ray burst lasting ∼ 50 s, with two major peaks of similar intensity spaced 9 s apart, and a third peak after 17 s. The fluxes of the three peaks are ∼ 4500, 5500, and 700 counts/s. The burst is in the field of Wide Field Camera (WFC) number 2 with a peak countrate of ∼ 300 counts/s. Taking into account the off-axis position, the peak flux corresponds to ∼ 4 Crab. From the WFC image, we have derived a first estimation of the position at α = 15^{h}28^{m}24^{s}, δ = +19^{o}40′.0 (equinox 2000.0); we conservatively estimate a 10′ error radius. The procedure for a TOO pointing of BeppoSAX Narrow Field Instruments has been started. Observations in all bands are solicited.”

PSR 0656+14

V. G. Kurt and B. V. Komberg, Astro Space Center, Russian Academy of Sciences; V. V. Sokolov and S. V. Zharikov, Special Astrophysical Observatory; and G. G. Pavlov, Pennsylvania State University, report: “The radio pulsar PSR 0656+14 was observed with the 6-m telescope on Nov. 11–12. Broadband BVRI CCD images (field-of-view 2′.4 × 2′.6; pixel size 0′′.274; seeing ∼ 1′′.5) show the following magnitudes of the pulsar: B = 24.85_{−0.16}^{+0.19}, V = 24.90_{−0.14}^{+0.16}, R = 24.52_{−0.11}^{+0.12}, I = 23.81_{−0.21}^{+0.27}. The opticalultraviolet spectrum, including the flux measured in the HST/FOC F130LP filter (cf. Pavlov, Stringfellow, and Cordova 1996, Ap.J. 467, 370) can be interpreted as a sum of the nonthermal power-law spectrum and thermal spectrum emitted by the surface of the neutron-star.”

1997 January 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green